Homepage › Solution manuals › Terence Tao › Probability Theory › Exercise 2.18 (Random variable is dependent of itself)

Exercise 2.18 (Random variable is dependent of itself)

Show that a random variable X taking values in a locally compact, σ-compact metric space is independent of itself (i.e. X and X are independent) if and only if X is almost surely equal to a constant.

Answers

In the following, let (Ω,F,P) denote the randomness source of X. The range R of X is σ-locally compact metric space, i.e.,

  • (σ-compact) R can be represented as a union ⋃ ⁡ n=1∞Cn of compact sets (Cn)n∈N;
  • (locally compact) every point r of R has a compact neighbourhood, i.e., there exists an open set U and a compact set K, such that r ∈ U ⊆ K;
  • (Hausdorff) every point r of R can be put into its own separate neighbourhood;

We will obviously use the Borel σ-algebra on R which also contains all compact sets and all singleton sets in our case. We now demonstrate both sides of the equivalence.

  • Suppose that X is independent of itself, i.e, for all S ∈B we have

    P(X ∈ S) = P(X ∈ S) ⋅ P(X ∈ S).

    Thus, from

    P(X ∈ S) = P(X ∈ S)2,

    we conclude that either P(X ∈ S) = 1 or that P(X ∈ S) = 0 for all S ∈B.
    Now suppose, for the sake of contradiction, that we have two values a,b in the image of X such that P(X = a) = c and P(X = b) = 1 − c for some 0 < c < 1. By the assumption of local compactness, we can find compact sets K,K′⊆ R such that a ∈ K and b ∈ K′. By monotonicity of the probability measure, this implies P(X ∈ K) ≥ c and P(X ∈ K′) ≥ 1 − c. But this leads to a contradiction with the previously derived fact claiming that any set has either a probability of zero or one because (1) we cannot have both P(X ∈ K) = 1 ≥ c and P(X ∈ K′) = 1 ≥ 1 − c; and (2) we cannot have one of theses probabilities to be zero as c > 0 and 1 − c > 0.

  • Suppose that there exists a ∈ R with P(X = a) = 1 and P(X≠a) = 0. For any set S ∈B we have two cases:

    • a ∈ S, in which case P(X ∈ S) ≥ P(X = a) and so 1 = P(X ∈ S) = P(X ∈ S)2.
    • a∉S, in which case P(X ∈ S) ≤ P(X∉a) and so 0 = P(X ∈ S) = P(X ∈ S)2.

    Hence, in any case we have P(X ∈ S) = P(X ∈ S) ⋅ P(X ∈ S).

User profile picture
2021-09-19 00:00
Comments
  • We can have both $1 \geq c$ and $1 \geq 1 - c$.
    isn • 2023-12-11